# CSTRL: Context-Driven Sequential Transfer Learning for Abstractive Radiology Report Summarization

### v1

@@ Abstract
A radiology report comprises several sections, including the Findings and Impression of the diagnosis. Automatically generating the Impression from the Findings is crucial for reducing radiologists’ workload and improving diagnostic accuracy. Pretrained models that excel in common abstractive summarization problems encounter challenges when applied to specialized medical domains largely due to the complex terminology and the necessity for accurate clinical context. Such tasks in medical domains demand extracting core information, avoiding context shifts, and maintaining proper flow. Misuse of medical terms can lead to drastic clinical errors. To address these issues, we introduce a sequential transfer learning that ensures key content extraction and coherent summarization. Sequential transfer learning often faces challenges like initial parameter decay and knowledge loss, which we resolve with the Fisher matrix regularization. Using MIMIC-CXR and Open-I datasets, our model, CSTRL—Context-driven Sequential TRansfer Learning—achieved state-of-the-art performance, showing 56.2% improvement in BLEU-1, 40.5% in BLEU-2, 84.3% in BLEU-3, 28.9% in ROUGE-1, 41.0% in ROUGE-2 and 26.5% in ROGUE-3 score over benchmark studies. We also analyze factual consistency scores while preserving the medical context. Our code is publicly available at TBA.
@@
Mst. Fahmida Sultana Naznin1, Adnan Ibney Faruq1, Mostafa Rifat Tazwar1, Md Jobayer2, Md. Mehedi Hasan Shawon2, Md Rakibul Hasan2,3, 1Bangladesh University of Engineering and Technology, Dhaka-1000, Bangladesh. 2BRAC University, Dhaka-1212, Bangladesh. 3Curtin University, Bentley WA 6102, Australia. Correspondence: rakibul.hasan@curtin.edu.au
@@ 1 Introduction
Writing Impression from Findings falls under abstractive summarization task that distills key clinical insights from diagnostic data Herts and Berland (2021). Extractive summarization selects sentences from the original text to create a summary, while abstractive summarization generates a new paragraph to summarize the content of the original document Cai et al. (2023). Abstractive summarization is more complex than extractive summarization but yields more flexible and concise summaries Cai et al. (2023). However, systematic evaluation of abstractive summarization across diverse domains is limited Zhang et al. (2020a). One challenge of abstractive summarization in the clinical domain is managing the balance between providing concise, clinically relevant information and avoiding excessive technical language. For example, “Liver metastases have enlarged” is more straightforward and actionable than “There is redemonstration of multiple liver lesions consistent with metastases, which have increased in size in the interval” Hartung et al. (2020). Including irrelevant Findings can create ambiguity, such as “diverticulosis without diverticulitis” Hartung et al. (2020). Diverticulosis indicates the presence of diverticula in the colon, which is typically benign. However, without proper context, it may raise concerns about potential complications and suggest that diverticulitis (inflammation of diverticula) could develop. To prioritize core sentences of Findings and preserve context, we propose a novel approach combining Sequential Transfer Learning with a Knowledge Distillation (KD) strategy, which automates the creation of Impression from the Findings of a radiology report.
@@
Deep learning techniques for abstractive summarization face specific challenges with radiology corpora. A key limitation is the difficulty in evaluating performance on datasets dense with radiology-specific terminology, along with significant differences in word distributions between general and radiology report domain Cai et al. (2023). Previous works (refer to Appendix A for details) have several shortcomings. Firstly, there is a lack of a proper methodology for generating Impression that focuses on primary observations Hartung et al. (2020). We propose an optimized Gap Sentence Generation (GSG) technique to identify key Findings. Secondly, Subtle differences in terminology can obscure the context and misrepresent the clinical scenario Codish and Shiffman (2005). In response, we propose a contextual tagging method. Thirdly, poor BLEU scores of recent studies indicate that the generated Impression section often fails to accurately reflect core points from the Findings section Cai et al. (2023). It affects the readability, informativeness, and reliability of reports for clinical use. Lastly, reducing dimensionality and computational complexity for real-time production is a challenge Meng et al. (2022); Che et al. (2015); Liu et al. (2020), for which we propose a KD method. Our key contributions in constructing Impression through proper clinical methodology with high performance and factual consistency are as follows:
@@
  • We introduce a sequential transfer learning framework with KD using the T5 (Text-to-Text Transfer Transformer) model, fine-tuned on a radiology corpus. This approach aims to transfer knowledge from the GSG task to summarization effectively gradually. To mitigate catastrophic forgetting during this sequential knowledge transfer, our model includes Fisher matrix regularization with penalty adjustment.
@@ 2 Method
We adopt a sequential two-step approach to fine-tune the T5 model Raffel et al. (2020) using a dataset on radiology corpus optimized by the GSG methodology Zhang et al. (2020a). In the first step, the model is trained on a GSG task, where it predicts masked sentences, thereby gaining a deeper understanding of the crucial statements for crafting coherent summaries in Impression section. In the second step, we fine-tune the model for summarization using the learned weights as initial parameters and employ Fisher matrix regularization for effective GSG-trained knowledge transfer to the summarization task. Then, the GSG fine-tuned model drives sequential knowledge distillation to achieve dimensionality reduction.
@@ 3.1 Dataset Description
We evaluate our model using two public chest X-ray report datasets: MIMIC-CXR Johnson et al. (2024) and Open-I National Library of Medicine (2024). MIMIC-CXR contains 227,835 reports, while Open-I, a public radiography dataset from Indiana University, originally included 7,430 reports. To ensure the quality of the data for medical report summarization, we apply several filtering criteria inspired by a GPT-Based Radiology Report Optimization study Ma et al. (2024). We remove (1) incomplete reports that lack either Finding or Impression, (2) reports with Finding section fewer than 10 words, and (3) reports with an impression section fewer than 2 words. We retain 121,975 reports from MIMIC-CXR and 3,312 reports from Open-I. We then merge these datasets, total of 125,287 reports. This merged dataset is divided into training, validation, and test sets using an 8:1:1 split. Refer to Table 8 in Appendix B for more information on the datasets.
@@ 3.2 Experimental Setup
Meta-Llama-3-1B 0.2 0.1 0.2 0.8 0.3 0.1
@@
We use PyTorch to implement our model on an NVIDIA A100 GPU. We configure the Teacher Model with 6 layers, 512 embedding dimensions, and 8 attention heads, while the Student Model has 3 layers, 128 dimensions, and 4 attention heads. Both use a vocabulary size of 32,128. Inputs are tokenized to 512 tokens, targets to 256. Models are trained for 20 epochs with a batch size of 32, using AdamW with a 0.003 learning rate. Logits are adjusted with a temperature of 20, and distillation loss is scaled by $\alpha$ = 0.7. ROUGE and BLEU with the T5-small are used for key token identification, masking 3 sentences per Findings.
@@ 3.4 Performance Evaluation and Ablation Study
We achieve remarkable performance in predicting masked sentences shown in Table 4 with ROUGE and BLEU values. In abstractive summarization, capturing inter-sentence relationships is crucial. We employ GSG, where sentences are masked based on ROUGE and BLUE, akin to an extractive summary. This method effectively accommodates both large and small datasets. For three batch sizes, the ROUGE and BLEU values are quite similar, with BLEU-1 at 55.56, BLEU-2 at 45.56, and BLEU-3 also at 45.56, showing consistent performance across various configurations.
@@
Table 4: Performance of GSG task with Different Hyperparameters
@@
We note that the performance metrics show identical values with and without the GSG framework, as shown in Table 2 for facing catastrophic forgetting. However, our experiments reveal differences beyond the third decimal point. Direct fine-tuning on the summarization task results in substantial updates to the model parameters, potentially compromising the knowledge gained during GSG training. After applying Fisher matrix regularization, we observe significant improvements in the results. These enhancements are due to the gradual adjustment of key parameters encapsulated within Fisher matrix. As a result, ROUGE-1, ROUGE-2, ROUGE-L, BLUE-1, BLUE-2 and BLUE-3 scores exhibit a notable 4.1%, 7.3%, 4.2%, 2.8%, 5.5%, 9.9% improvement respectively. We explore an alternative to Fisher matrix regularization by implementing a gradual layer unfreezing strategy. Initially, we freeze certain layers to retain knowledge from the previous fine-tuned task (GSG). Then, we slowly unfreeze these layers to minimize the risk of losing critical information. This approach reduces performance, decreasing ROUGE-1, ROUGE-2, ROUGE-L, BLEU-1, BLEU-2, and BLEU-3 scores by 4.5%, 4.6%, 4.2%, 2.7%, 6.8%, and 8.7%, respectively. (In Table 2, we test GSG with layer unfreezing.).
Model Readability Factual Correct. Informativeness Redundancy Completeness
@@
Table 5: Human evaluation results based on summarization quality.
@@
As shown in Table 3, we scale the network parameters to investigate the CSTRL Student model’s (CSTRL-S) performance compared to the CSTRL Teacher model (CSTRL-T). Reducing the parameter size significantly drops the performance from 20.8% to 15.3%, 28.04% to 21.9%, 20.5% to 13.6%, 13.2% to 6%, 27.8% to 22.5%, and 37.5% to 32.39% in ROUGE-1, ROUGE-2, ROUGE-L, BLEU-1, BLEU-2, and BLEU-3, respectively. In Table 6, we measure complexity and computational efficiency. Inference time is reduced by up to 63.1%.
@@ 3.6 Human Evaluation
CSTRL-T 71.7 35.8 60.5 0.521s
CSTRL-S (×$\times 8$) 17.9 9.0 11.2 0.330s
CSTRL-S (×$\times 16$) 10.6 5.3 6.4 0.220s
CSTRL-S (×$\times 32$) 5.6 2.8 3.63 0.192s
@@ 4 Conclusion
The study is conducted on two datasets, both containing predominantly short sentences in the Impression section, with an average length of 5.52 words. A key limitation is that the model’s predicted impressions are similarly concise, likely due to the bias in the training data. As a result, the model may encounter difficulties in generating longer, more detailed impressions.
@@ Acknowledgement
This work was supported by resources provided by the Pawsey Supercomputing Research Centre’s Setonix Supercomputer (https://doi.org/10.48569/18sb-8s43), with funding from the Australian Government and the Government of Western Australia. We also gratefully acknowledge the valuable contributions of the five volunteers who dedicated their time and expertise to evaluate our model.
Appendix A Related Work
In recent years, deep learning has significantly improved neural abstractive summarization tasks in NLP Rush et al. (2015); Chopra et al. (2016); Tan et al. (2017); Zhou et al. (2017); Li et al. (2017). Traditional models are primarily trained on general datasets like CNN/Daily Mail Nallapati et al. (2016) and Gigaword Corpus Sutskever et al. (2014). The attention-based seq2seq model by Rush et al. (2015) and its extension by Chopra et al. (2016) laid the groundwork for these advancements. Recently, large-scale pre-trained language models have shown impressive summarization results Liu and Chen (2019); Karn et al. (2021); Raffel et al. (2020); Kieuvongngam et al. (2020); Chang et al. (2021). Kieuvongngam et al. (2020) employed BERT and GPT for summarizing COVID-19 research. However, unique word distributions in radiology corpora limit the applicability of these techniques Cai et al. (2023).
Zhang et al. (2018) first explored automatic radiology Impression generation, followed by MacAvaney et al. (2019), who introduced an ontology-aware pointer-generator model to enhance summarization quality. A background-augmented pointer-generator network with copy and background-guided decoding was proposed Zhang et al. (2020b), and a word graph captured critical words and relations Hu et al. (2021). Anatomies were extracted, radiographs encoded, and fused with anatomy-enhanced co-attention Hu et al. (2023). Sotudeh Gharebagh et al. (2020) enhanced clinical summarization by augmenting ontological terms, and Hu et al. (2022) integrated findings by a unified framework with knowledge via text and graph encoders. Karn et al. (2022) proposed a two-step extractive-abstractive method using a Differentiable Multi-Agent Actor-Critic framework.
A hybrid retrieval-generation agent Li et al. (2018) integrated human knowledge with neural networks for medical report generation. Models like OpenAI GPT Chang et al. (2021), BERT Devlin et al. (2019), ELMo Peters et al. (2018), and XLNet Yang et al. (2019) have improved various NLP tasks through extensive external knowledge. BioBERT Lee et al. (2020) captures semantic features in the biomedical field, pre-trained on large corpora like PubMed abstracts and PMC articles. Recently, Cai et al. (2023) proposed a pre-trained language model, ChestXrayBert, specifically designed for summarizing chest radiology reports. However, these studies may struggle with effective Impression generation methodologies.
Dataset # Reports Findings (Avg.) Impressions (Avg.)
# S # W/S # Source W # S # W/S # Source W
MIMIC-CXR 121,975 5.47 10.08 55.09 1.94 8.50 16.46
Open-I 3,312 4.62 8.03 37.06 1.81 5.52 9.98
  • S: Sentence, W: Words
Table 8: Summary of dataset statistics of Findings and Impression
Appendix B Additional Details of the Datasets
Summary statistics of both datasets are shown in Table 8. MIMIC-CXR comprises a significantly larger collection of reports, with a total of 121,975 entries, while Open-I contains 3,312 reports. Using two datasets enhances the model’s robustness and generalizability, allowing it to learn diverse patterns and improve performance across various scenarios. When analyzing Findings, MIMIC-CXR exhibits an average of 5.47 sentences per finding, with each sentence containing approximately 10.08 words. It has an average of 55.09 words per Findings, highlighting the dataset’s richness in detail and complexity. Conversely, Open-I has slightly fewer sentences per finding, averaging 4.62 sentences, with each sentence comprising 8.03 words and a total of 37.06 words.
Impression reveals similar trends. MIMIC-CXR has an average of 1.94 sentences per impression, with an average of 8.50 words per sentence, leading to an average of 16.46 words. In contrast, Open-I averages 1.81 sentences per impression, with 5.52 words per sentence and 9.98 source words.
Appendix C Hyperparameter Tuning
We conduct a grid search throughout all experiments for tuning hyperparameters. Table 4 shows a summary of the optimal performance achieved through GSG fine-tuning for each batch size. The table shows how varying these hyperparameters impacts the evaluation metrics: ROUGE-1 (R-1), ROUGE-2 (R-2), ROUGE-L (R-L), BLEU-1 (B-1), BLEU-2 (B-2), and BLEU-3 (B-3). As the batch size increases from 8 to 32, with a corresponding adjustment in learning rate, there is a noticeable improvement in the scores for ROUGE and BLEU metrics. For instance, the highest ROUGE-1 score of 39.7 is achieved with a batch size of 32 and a learning rate of 0.003, indicating that fine-tuning these hyperparameters can enhance the model’s summarization performance.

### v2

@@ Abstract
A radiology report comprises several sections, including the Findings and Impression of the diagnosis. Automatically generating the Impression from the Findings is crucial for reducing radiologists’ workload and improving diagnostic accuracy. Pretrained models that excel in common abstractive summarization problems encounter challenges when applied to specialized medical domains, largely due to the complex terminology and the necessity for accurate clinical context. Such tasks in medical domains demand extracting core information, avoiding context shifts, and maintaining proper flow. Misuse of medical terms can lead to drastic clinical errors. To address these issues, we introduce a sequential transfer learning that ensures key content extraction and coherent summarization. Sequential transfer learning often faces challenges like initial parameter decay and knowledge loss, which we resolve with the Fisher matrix regularization. Using MIMIC-CXR and Open-I datasets, our model, CSTRLContext-driven Sequential TRansfer Learning—achieved state-of-the-art performance: 56.2% improvement in BLEU-1, 40.5% in BLEU-2, 84.3% in BLEU-3, 28.9% in ROUGE-1, 41.0% in ROUGE-2 and 26.5% in ROGUE-3 score over benchmark methods. We further analyze factual consistency scores while preserving the medical context. Our code is publicly available at https://github.com/fahmidahossain/Report_Summarization.
@@
Mst. Fahmida Sultana Naznin1, Adnan Ibney Faruq1, Mostafa Rifat Tazwar1, Md Jobayer2, Md. Mehedi Hasan Shawon2, Md Rakibul Hasan2,3 1Bangladesh University of Engineering and Technology, Dhaka 1000, Bangladesh 2BRAC University, Dhaka 1212, Bangladesh 3Curtin University, Bentley WA 6102, Australia Correspondence: rakibul.hasan@curtin.edu.au
@@ 1 Introduction
Writing Impression from Findings falls under the abstractive summarization task that distills key clinical insights from diagnostic data Herts and Berland (2021). Extractive summarization selects sentences from the original text to create a summary, while abstractive summarization generates a new paragraph to summarize the content of the original document Cai et al. (2023). Abstractive summarization is more complex than extractive summarization but yields more flexible and concise summaries Cai et al. (2023). However, systematic evaluation of abstractive summarization across diverse domains is limited Zhang et al. (2020a). One challenge of abstractive summarization in the clinical domain is managing the balance between providing concise, clinically relevant information and avoiding excessive technical language. For example, “Liver metastases have enlarged” is more straightforward and actionable than “There is redemonstration of multiple liver lesions consistent with metastases, which have increased in size in the interval” Hartung et al. (2020). Including irrelevant Findings can create ambiguity. For example, “diverticulosis without diverticulitis” Hartung et al. (2020): diverticulosis indicates the presence of diverticula in the colon, which is typically benign, but, without proper context, it may raise concerns about potential complications and suggest that diverticulitis (inflammation of diverticula) could develop. To prioritize core sentences of Findings and preserve context, we propose a novel approach combining Sequential Transfer Learning with a Knowledge Distillation (KD) strategy, which automates the creation of Impression from the Findings of a radiology report.
@@
Deep learning techniques for abstractive summarization face specific challenges with radiology corpora. A key limitation is the difficulty in evaluating performance on datasets dense with radiology-specific terminology, along with significant differences in word distributions between general and radiology report domains Cai et al. (2023). Previous works (Section 2) have several shortcomings. Firstly, there is a lack of a proper methodology for generating Impression that focuses on primary observations Hartung et al. (2020). We propose an optimized Gap Sentence Generation (GSG) technique to identify key Findings. Secondly, subtle differences in terminology can obscure the context and misrepresent the clinical scenario Codish and Shiffman (2005). In response, we propose a contextual tagging method. Thirdly, poor BLEU scores of recent studies indicate that the generated Impression section often fails to accurately reflect core points from the Findings section Cai et al. (2023). It affects the readability, informativeness, and reliability of reports for clinical use. Lastly, reducing dimensionality and computational complexity for real-time production is a challenge Meng et al. (2022); Che et al. (2015); Liu et al. (2020), for which we propose a KD method. Our key contributions in constructing Impression through proper clinical methodology with high performance and factual consistency are as follows:
@@
  • We introduce a sequential transfer learning framework with KD using the T5 (Text-to-Text Transfer Transformer) model, fine-tuned on a radiology corpus. This approach aims to transfer knowledge from the GSG task to summarization effectively and gradually. To mitigate catastrophic forgetting during this sequential knowledge transfer, our model includes Fisher matrix regularization with penalty adjustment.
@@
  • We compare our proposed summarization model, CSTRL, against state-of-the-art models trained on the same dataset. We further train decoder-only models for comparison. CSTRL consistently outperforms all baselines across various settings. We further analyze factual consistency and conduct a comprehensive human evaluation to validate the quality and reliability of the generated summaries.
@@
2 Related Work
In recent years, deep learning has significantly improved neural abstractive summarization tasks in NLP Rush et al. (2015); Chopra et al. (2016); Tan et al. (2017); Zhou et al. (2017); Li et al. (2017). Traditional models are primarily trained on general datasets like CNN/Daily Mail Nallapati et al. (2016) and Gigaword Corpus Sutskever et al. (2014). The attention-based seq2seq model by Rush et al. (2015) and its extension by Chopra et al. (2016) laid the groundwork for these advancements. Recently, large-scale pre-trained language models have shown impressive summarization results Karn et al. (2021); Raffel et al. (2020); Kieuvongngam et al. (2020); Chang et al. (2021). Kieuvongngam et al. (2020) employed BERT and GPT for summarizing COVID-19 research. However, unique word distributions in radiology corpora limit the applicability of these techniques Cai et al. (2023).
Zhang et al. (2018) first explored automatic radiology Impression generation, followed by MacAvaney et al. (2019), who introduced an ontology-aware pointer-generator model to enhance summarization quality. A background-augmented pointer-generator network with copy and background-guided decoding was proposed in Zhang et al. (2020b), and a word graph captured critical words and relations in Hu et al. (2021). Anatomies were extracted, radiographs encoded, and fused with anatomy-enhanced co-attention in Hu et al. (2023). Sotudeh Gharebagh et al. (2020) enhanced clinical summarization by augmenting ontological terms, and Hu et al. (2022) integrated Findings by a unified framework with knowledge via text and graph encoders. Karn et al. (2022) proposed a two-step extractive-abstractive method using a Differentiable Multi-Agent Actor-Critic framework.
Li et al. (2018) proposed a hybrid retrieval-generation agent that integrates human knowledge with neural networks for medical report generation. Models like OpenAI GPT Chang et al. (2021), BERT Devlin et al. (2019), ELMo Peters et al. (2018), and XLNet Yang et al. (2019) have improved various NLP tasks through extensive external knowledge. BioBERT Lee et al. (2020) captures semantic features in the biomedical field, pre-trained on large corpora like PubMed abstracts and PMC articles. Recently, Cai et al. (2023) proposed a pre-trained language model, ChestXrayBert, specifically designed for summarizing chest radiology reports. However, these studies may struggle with effective Impression generation methodologies.
@@ 3 Method
We adopt a sequential two-step approach to fine-tune the T5 model Raffel et al. (2020) using a dataset on radiology corpus optimized by the GSG methodology Zhang et al. (2020a). In the first step, the model is trained on a GSG task, where it predicts masked sentences, thereby gaining a deeper understanding of the crucial statements for crafting coherent summaries in the Impression section. In the second step, we fine-tune the model for summarization using the learned weights as initial parameters and employ Fisher matrix regularization for effective GSG-trained knowledge transfer to the summarization task. Then, the GSG fine-tuned model drives sequential knowledge distillation to achieve dimensionality reduction.
@@ 4 Experiment and Evaluation
Dataset # Reports Findings (Avg.) Impressions (Avg.)
# S # W/S # Source W # S # W/S # Source W
MIMIC-CXR 121,975 5.47 10.08 55.09 1.94 8.50 16.46
Open-I 3,312 4.62 8.03 37.06 1.81 5.52 9.98
  • S: Sentence, W: Words
Table 1: Summary of dataset statistics of Findings and Impression
@@ 4.1 Dataset Description
We evaluate our model using two public chest X-ray report datasets: MIMIC-CXR Johnson et al. (2024) and Open-I National Library of Medicine (2024), as shown in Table 1. MIMIC-CXR contains 227,835 reports, while Open-I, a public radiography dataset from Indiana University, originally included 7,430 reports. Findings in the MIMIC-CXR dataset exhibit an average of 5.47 sentences per Findings, with each sentence containing approximately 10.08 words. It has an average of 55.09 words per Findings, which suggests this dataset is rich in detail and complexity. Conversely, Open-I has slightly fewer sentences per Findings, averaging 4.62 sentences, with each sentence comprising 8.03 words and a total of 37.06 words. Impression reveals similar trends. MIMIC-CXR has an average of 1.94 sentences per Impression, with an average of 8.50 words per sentence, leading to an average of 16.46 words. In contrast, Open-I averages 1.81 sentences per Impression, with 5.52 words per sentence and 9.98 source words.
To ensure the quality of the data for medical report summarization, we apply several filtering criteria inspired by a GPT-Based Radiology Report Optimization study Ma et al. (2024). We remove (1) incomplete reports that lack either Findings or Impression, (2) reports with a Findings section of fewer than 10 words, and (3) reports with an Impression section of fewer than 2 words. We retain 121,975 reports from MIMIC-CXR and 3,312 reports from Open-I. We then merge these datasets, a total of 125,287 reports. This merged dataset is divided into training, validation, and test sets using an 8:1:1 split.
@@ 4.2 Experimental Setup
Meta-Llama-3-1B (Zero-shot) 0.2 0.1 0.2 0.8 0.3 0.1
Meta-Llama-3-1B (Few-shot) 30.5 12.9 27.9 31.2 55.2 67.4
@@
We use PyTorch to implement our model on an NVIDIA A100 GPU. We configure the Teacher Model with 6 layers, 512 embedding dimensions, and 8 attention heads, while the Student Model has 3 layers, 128 dimensions, and 4 attention heads. Both use a vocabulary size of 32,128. Inputs are tokenized to 512 tokens, targets to 256. Models are trained for 20 epochs with a batch size of 32, using AdamW with a 0.003 learning rate. Logits are adjusted with a temperature of 20, and distillation loss is scaled by $\alpha$ = 0.7. ROUGE and BLEU with the T5-small are used for key token identification, masking 3 sentences per Findings. We conduct a grid search throughout all experiments for tuning hyperparameters.
@@ 4.4 Performance Evaluation and Ablation Study
As shown in Table 5, we predict masked sentences and evaluate using ROUGE and BLEU values for different batch sizes. In abstractive summarization, capturing inter-sentence relationships is crucial. We employ GSG, where sentences are masked based on ROUGE and BLUE, akin to an extractive summary. This method effectively accommodates both large and small datasets. Across the three batch sizes, the ROUGE and BLEU scores are quite similar, showing consistent performance across various configurations. The best configuration appears to be a batch size of 32 and a learning rate of 0.003, which provides a BLEU-1 of 55.7, a BLEU-2 of 35.8, and a BLEU-3 of 28.1.
@@
Table 5: Performance of GSG task with Different Hyperparameters.
@@
We note that the performance metrics show identical values with and without the GSG framework, as shown in Table 3, for facing catastrophic forgetting. However, our experiments reveal differences beyond the third decimal point. Direct fine-tuning on the summarization task results in substantial updates to the model parameters, potentially compromising the knowledge gained during GSG training. After applying Fisher matrix regularization, we observe significant improvements in the results. These enhancements are due to the gradual adjustment of key parameters encapsulated within the Fisher matrix. As a result, ROUGE-1, ROUGE-2, ROUGE-L, BLUE-1, BLUE-2 and BLUE-3 scores exhibit a notable 4.1%, 7.3%, 4.2%, 2.8%, 5.5% and 9.9% improvements, respectively. We explore an alternative to Fisher matrix regularization by implementing a gradual layer unfreezing strategy. Initially, we freeze certain layers to retain knowledge from the previous fine-tuned task (GSG). Then, we slowly unfreeze these layers to minimize the risk of losing critical information. This approach reduces performance, decreasing ROUGE-1, ROUGE-2, ROUGE-L, BLEU-1, BLEU-2, and BLEU-3 scores by 4.5%, 4.6%, 4.2%, 2.7%, 6.8%, and 8.7%, respectively (Table 3).
Model Readability Factual Correctness Informativeness Redundancy Completeness
@@
Table 6: Human evaluation results based on summarization quality. Higher scores indicate better performance.
@@
As shown in Table 4, we scale the network parameters to investigate the CSTRL Student model’s (CSTRL-S) performance compared to the CSTRL Teacher model (CSTRL-T). Reducing the parameter size significantly drops the performance from 20.8% to 15.3%, 28.04% to 21.9%, 20.5% to 13.6%, 13.2% to 6%, 27.8% to 22.5%, and 37.5% to 32.39% in ROUGE-1, ROUGE-2, ROUGE-L, BLEU-1, BLEU-2, and BLEU-3, respectively. In Table 7, we measure complexity and computational efficiency. Inference time is reduced by up to 63.1% compared to CSTRL-T. We compare our model, CSTRL, with different settings against widely used baselines. We observe that CSTRL models (T and S variants) are highly efficient. They outperform existing models by reducing computational demands (GFLOPS, GMACs) while maintaining compact parameter sizes and competitive performance. The CSTRL-S variants (×8, ×16, ×32) exhibit a scalable efficiency trade-off: as the model scales down (8× to 32×), computational load decreases, achieving 2.5 to 8 times faster inference than ClinicalBERT or BART and 7 to 20 times faster than XrayGPT, despite using 10 to 100 times fewer parameters. In contrast, larger models like Meta-Llama-3-1B and XrayGPT demand significantly higher computational resources—XrayGPT requires 45 times more GMACs than CSTRL-S 32×—and have considerably slower inference (4s vs. 0.19s). We achieve near-real-time performance with minimal resources, bridging lightweight deployment and clinical-grade AI.
@@ 4.6 Human Evaluation
CSTRL-T 71.7 35.8 60.50M 0.521s
CSTRL-S (×$\times 8$) 17.9 9.0 11.20M 0.330s
CSTRL-S (×$\times 16$) 10.6 5.3 6.40M 0.220s
CSTRL-S (×$\times 32$) 5.6 2.8 3.63M 0.192s
Meta-Llama-3-1B 1270 632.7 1.24B 1.527s
ClinicalBERT Huang et al. (2020) 50 40 110M 1s
BART Lewis et al. (2019) 100 200 140M 2s
XrayGPT Thawakar et al. (2024) 250 500 350M 4s
@@ 5 Conclusion
Our proposed model is evaluated on two datasets, both containing predominantly short sentences in the Impression section. A key limitation is that the model’s predicted Impressions are similarly concise, likely due to the bias in the training data. As a result, the model may encounter difficulties in generating longer, more detailed impressions.
@@ Acknowledgement
This work was supported by resources provided by the Pawsey Supercomputing Research Centre with funding from the Australian Government and the Government of Western Australia. We also gratefully acknowledge the valuable contributions of the five volunteers who dedicated their time and expertise to evaluate our model outputs.